# Unleashing the Denoising Capability of Diffusion Prior for Solving Inverse Problems

### v1

@@ Abstract
The recent emergence of diffusion models has significantly advanced the precision of learnable priors, presenting innovative avenues for addressing inverse problems. Since inverse problems inherently entail maximum a posteriori estimation, previous works have endeavored to integrate diffusion priors into the optimization frameworks. However, prevailing optimization-based inverse algorithms primarily exploit the prior information within the diffusion models while neglecting their denoising capability. To bridge this gap, this work leverages the diffusion process to reframe noisy inverse problems as a two-variable constrained optimization task by introducing an auxiliary optimization variable. By employing gradient truncation, the projection gradient descent method is efficiently utilized to solve the corresponding optimization problem. The proposed algorithm, termed ProjDiff, effectively harnesses the prior information and the denoising capability of a pre-trained diffusion model within the optimization framework. Extensive experiments on the image restoration tasks and source separation and partial generation tasks demonstrate that ProjDiff exhibits superior performance across various linear and nonlinear inverse problems, highlighting its potential for practical applications. Code is available at https://github.com/weigerzan/ProjDiff/.
@@ 1 Introduction
Denoising diffusion models have achieved tremendous success in the field of generative modeling [1, 2, 3, 4, 5]. Their remarkable ability to capture data priors enables effective guidance for solving inverse problems [6], which are widely exploited in image restoration [7, 8, 9, 10, 11, 12], medical image processing [13, 14], 3D vision [15], audio processing [16, 17] and beyond.
@@
To fully utilize both the prior information and the denoising capability inherent in diffusion models within the optimization framework, we introduce an auxiliary optimization variable to accommodate the influence of the observation noise. We derive a novel two-variable objective based on the properties of the diffusion process and transform the inverse problem into a constrained optimization task. Through gradient truncation, we obtain an approximation of the stochastic gradient of the objective, effectively sidestepping significant computational overhead. The proposed algorithm, termed ProjDiff, tackles the inverse problem by employing the concept of projection gradient descent to solve the corresponding optimization task. We also discuss the noise-free version of ProjDiff as a special case, which, compared to other optimization-based methods, ensures superior consistency between the generated results and observations. ProjDiff’s applicability also extends beyond linear observations to encompass nonlinear functions, which enhances its competitiveness and versatility.
@@ 2.1 Denoising diffusion models
Here $\bm{\bm{\mu}_{\bm{\theta}}}$ serves as an estimation of $x_{0}$ at time $t$, typically implemented through a parameterized neural network referred to as a “$\bm{\mu}$-predictor” [^2] The case of $\bm{\epsilon}$-predictor is similar, with the conversion given by $\bm{\bm{\mu}}_{\bm{\theta}}(x_{t})=\left(x_{t}-\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}_{\bm{\theta}}(x_{t})\right)/\sqrt{\overline{\alpha}_{t}}$.. The variance $\sigma_{t}^{2}\in[0,1-\overline{\alpha}_{t-1}]$ can take various values. When $\sigma_{t}=\sqrt{\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}}\sqrt{1-\frac{\overline{\alpha}_{t}}{\overline{\alpha}_{t-1}}}$, the reverse process aligns with DDPM [2]; while when retaining $\sigma_{t}=\tilde{\sigma}_{t}$ as an adjustable parameter, one obtains DDIM [3].
@@ 3.4 Time steps, initialization, and Restricted Encoding
The complete implementation of ProjDiff for noisy observations is outlined in Algorithm 1. More variations of ProjDiff used in our experiments can be found in Appendix G.
@@ 4.1 Image restoration
Linear tasks. We demonstrate the performance of ProjDiff across three linear image restoration tasks on ImageNet [32]: 4×4$\times$4 × super-resolution (with average pooling), $50$% random inpainting, and Gaussian deblurring. Comparisons are conducted against prominent diffusion-based image restoration algorithms in recent years, including DDRM [20], DDNM [21], DPS [23], and RED-diff [24]. The performance of the least square solution ($\hat{x}_{0}=\mathbf{A}^{\dagger}y$) is reported as a baseline. For a fair comparison, DDRM, DDNM, RED-diff, and ProjDiff utilize $100$ function evaluations, while DPS uses $1000$ function evaluations as its performance significantly declines with $100$ steps. The pre-trained model on ImageNet is sourced from [33], specifically the $256\times 256$ model without classifier guidance. Testing is conducted on a 1k sub-testset of ImageNet consistent with [21]. Metrics include PSNR, SSIM [34], LPIPS [35], and FID [36], with all LPIPS values multiplied by $100$ for clarity.
@@
Nonlinear tasks. We evaluate the effectiveness of ProjDiff for nonlinear observations through the challenging phase retrieval task on the FFHQ dataset [38]. We compare ProjDiff against DPS [23] and RED-diff [24] as they are applicable to nonlinear observations. Table 2 presents the results for both noise-free and noisy scenarios. Given the inherently ill-posedness of phase retrieval, ProjDiff requires more steps to achieve superior results. We set the number of function evaluations to $1000$ for all three algorithms. The results demonstrate that ProjDiff excels in addressing nonlinear inverse problems, regardless of whether the observation is noise-free or noisy. Some visualization results are shown in Figure 3. Further experiments on the high dynamic range (HDR) task can be found in Appendix C.
@@ 5 Conclusion
In the appendix, we provide further details and additional experimental results omitted in the main text. First, we present the complete derivations of the propositions for the ProjDiff algorithm (Appendix A), and then provide the derivations for ProjDiff in the variance exploding setting (Appendix B). Next, we present additional numerical results. For the image restoration tasks, we report the results of noise-free scenarios on the ImageNet dataset, both noise-free and noisy scenarios on the CelebA dataset, the nonlinear HDR task on the FFHQ dataset, and the ablation studies (Appendix C). For source separation and partial generation tasks, results on the MSDM model, and ablation experiments with waveform visualizations are shown in Appendix D. Then we provide the details of the experiments, including the mathematical form of the inverse problems, and the details for calculating metrics (Appendix E). The implementation details of ProjDiff for various inverse problems are discussed in Appendix F, especially the details for the phase retrieval task and high dynamic range task. Algorithm blocks for all kinds of ProjDiff used in this work are outlined in Appendix G. Finally, we include more visualizations of the restoration results to present the performance of ProjDiff intuitively in Appendix H.
@@ Appendix A Derivation for ProjDiff in VP diffusion
When $x>t_{a}$ the process is Markov and thus the conditional distribution is independent of $x_{0}$, while when $x\leq t_{a}$ the conditional mean of $x_{t}$ is generally a linear combination of $x_{0}$ and $x_{t_{a}}$.
@@
Then based on these preliminaries, we derive the ELBO of $\log p_{\bm{\theta}}(x_{0},x_{t_{a}})$ as follows:
@@
where the coefficient $g(t)$ is omitted as it can be scaled into the step size. Here we reach the result of Proposition 2 in the main text. Then we truncate the gradient through $\bm{\bm{\mu}_{\bm{\theta}}}$, i.e., we let
@@ C.1 Additional results
Here, we present the omitted experiments for image restoration tasks. The results for noise-free restoration tasks on ImageNet are shown in Table 5, the results for noise-free and noisy restoration tasks on CelebA are shown in Table 6 and 7, respectively, and the results for nonlinear high dynamic range (HDR) task are in Table 8. ProjDiff demonstrates highly competitive performance on all these tasks.
@@ E.2 Hyperparameters for the compared algorithms
DDNM. DDNM uses the recommended parameter $\eta=0.85$ from [21].
DPS. Following the settings in [23], the step size is set to $\xi_{t}=\xi/\left|\left|y-\mathcal{A}(\bm{\bm{\mu}_{\bm{\theta}}}(x_{t}))\right|\right|_{2}$. On ImageNet, $\xi$ is set to $1$ for super-resolution and inpainting tasks, and $0.4$ for Gaussian deblurring. Since [23] did not conduct experiments on CelebA, we use the parameters on FFHQ for the CelebA dataset as both FFHQ and CelebA are datasets of faces, i.e. $\xi=1$ for super-resolution, inpainting, and Gaussian deblurring. For phase retrieval and high dynamic range tasks, $\xi=0.4$.
@@ F.3 Details for nonlinear observations
Appendix G Algorithm blocks

### v2

@@ Abstract
The recent emergence of diffusion models has significantly advanced the precision of learnable priors, presenting innovative avenues for addressing inverse problems. Previous works have endeavored to integrate diffusion priors into the maximum a posteriori estimation (MAP) framework and design optimization methods to solve the inverse problem. However, prevailing optimization-based rithms primarily exploit the prior information within the diffusion models while neglecting their denoising capability. To bridge this gap, this work leverages the diffusion process to reframe noisy inverse problems as a two-variable constrained optimization task by introducing an auxiliary optimization variable that represents a ’noisy’ sample at an equivalent denoising step. The projection gradient descent method is efficiently utilized to solve the corresponding optimization problem by truncating the gradient through the $\mu$-predictor. The proposed algorithm, termed ProjDiff, effectively harnesses the prior information and the denoising capability of a pre-trained diffusion model within the optimization framework. Extensive experiments on the image restoration tasks and source separation and partial generation tasks demonstrate that ProjDiff exhibits superior performance across various linear and nonlinear inverse problems, highlighting its potential for practical applications. Code is available at https://github.com/weigerzan/ProjDiff/.
@@ 1 Introduction
Denoising diffusion models have achieved tremendous success in the field of generative modeling [1, 2, 3, 4, 5]. Their remarkable ability to capture data priors provides promising avenues for solving inverse problems [6], which are widely exploited in image restoration [7, 8, 9, 10, 11, 12], medical image processing [13, 14], 3D vision [15], audio processing [16, 17] and beyond.
@@
To fully utilize both the prior information and the denoising capability inherent in diffusion models within the optimization framework, we introduce an auxiliary optimization variable to accommodate the influence of the observation noise. We derive a novel two-variable objective based on the properties of the diffusion process and transform the inverse problem into a constrained optimization task. Through gradient truncation, we obtain a more practical approximation of the stochastic gradient of the objective which sidesteps significant computational overhead. The proposed algorithm, termed ProjDiff, tackles the inverse problem by employing the concept of projection gradient descent to solve the corresponding optimization task. We also discuss the noise-free version of ProjDiff as a special case, which, compared to other optimization-based methods, ensures superior consistency between the generated results and observations. ProjDiff’s applicability also extends beyond linear observations to encompass nonlinear functions, which enhances its competitiveness and versatility.
@@ 2.1 Denoising diffusion models
Here $\bm{\bm{\mu}_{\bm{\theta}}}$ serves as an estimation of $x_{0}$ at time $t$, typically implemented through a parameterized neural network referred to as a “$\bm{\mu}$-predictor” [^2] The case of $\bm{\epsilon}$-predictor is similar, with the conversion given by $\bm{\bm{\mu}}_{\bm{\theta}}(x_{t})=\left(x_{t}-\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon}_{\bm{\theta}}(x_{t})\right)/\sqrt{\overline{\alpha}_{t}}$.. The variance $\sigma_{t}^{2}\in[0,1-\overline{\alpha}_{t-1}]$ can take various values. When $\sigma_{t}=\sqrt{\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}}\sqrt{1-\frac{\overline{\alpha}_{t}}{\overline{\alpha}_{t-1}}}$, the reverse process aligns with DDPM [2]; while when retaining $\sigma_{t}=\tilde{\sigma}_{t}$ as an adjustable parameter, one obtains DDIM [3]. We refer to [19, 29, 30, 31] for more theoretical guarantees of diffusion models.
@@ 3.4 Time steps, initialization, and Restricted Encoding
The complete implementation of ProjDiff for noisy observations is outlined in Algorithm 1. More variations of ProjDiff used in our experiments can be found in Appendix H.
@@ 4.1 Image restoration
Linear tasks. We demonstrate the performance of ProjDiff across three linear image restoration tasks on ImageNet [35]: 4×4$\times$4 × super-resolution (with average pooling), $50$% random inpainting, and Gaussian deblurring. Comparisons are conducted against prominent diffusion-based image restoration algorithms in recent years, including DDRM [20], DDNM+ [21], DPS [23], and RED-diff [24]. The performance of the least square solution ($\hat{x}_{0}=\mathbf{A}^{\dagger}y$) is reported as a baseline. For a fair comparison, DDRM, DDNM, RED-diff, and ProjDiff utilize $100$ function evaluations, while DPS uses $1000$ function evaluations as its performance significantly declines with $100$ steps. The pre-trained model on ImageNet is sourced from [36], specifically the $256\times 256$ model without classifier guidance. Testing is conducted on a 1k sub-testset of ImageNet consistent with [21]. Metrics include PSNR, SSIM [37], LPIPS [38], and FID [39], with all LPIPS values multiplied by $100$ for clarity.
@@
Nonlinear tasks. We evaluate the effectiveness of ProjDiff for nonlinear observations through the challenging phase retrieval task on the FFHQ dataset [41]. We compare ProjDiff against DPS [23] and RED-diff [24] as they are applicable to nonlinear observations. Traditional baseline algorithms are also reported, including the Error Reduction (ER) algorithm [42], Hybrid Input-Output (HIO) [43], and Oversampling Smoothness (OSS) [44]. Table 2 presents the results for both noise-free and noisy scenarios. One sample is generated per image from each algorithm. Given the inherently ill-posedness of phase retrieval, ProjDiff requires more steps to achieve superior results. We set the number of function evaluations to $1000$ for all three algorithms. The results demonstrate that ProjDiff excels in addressing nonlinear inverse problems, regardless of whether the observation is noise-free or noisy. Some visualization results are shown in Figure 3. Further experiments on the high dynamic range (HDR) task can be found in Appendix C.
@@ 5 Conclusion
Acknowledgement
This work was supported by NSAF (Grant No. U2230201) and a Grant from the Guoqiang Institute, Tsinghua University.
@@ Acknowledgement
In the appendix, we provide further details and additional experimental results omitted in the main text. First, we present the complete derivations of the propositions for the ProjDiff algorithm (Appendix A), and then provide the derivations for ProjDiff in the variance exploding setting (Appendix B). Next, we present additional numerical results. For the image restoration tasks, we report the results of noise-free scenarios on the ImageNet dataset, both noise-free and noisy scenarios on the CelebA dataset, the nonlinear HDR task on the FFHQ dataset, and the ablation studies (Appendix C). For source separation and partial generation tasks, results on the MSDM model, and ablation experiments with waveform visualizations are shown in Appendix D. Then we provide the details of the experiments, including the mathematical form of the inverse problems, and the details for calculating metrics (Appendix E). The implementation details of ProjDiff for various inverse problems are discussed in Appendix F, especially the details for the phase retrieval task and high dynamic range task. Discussion about the performance gap between ProjDiff and the MAP framework is presented in Appendix G. Algorithm blocks for all kinds of ProjDiff used in this work are outlined in Appendix H. Finally, we include more visualizations of the restoration results to present the performance of ProjDiff intuitively in Appendix I.
@@ A.1 Preliminaries and lemmas
Proof.
When $t>t_{a}$, (30) holds as the chain between $t_{a}$ and $T$ Markov.
When $t<t_{a}$, (30) and (31) holds when $t=t_{a}$ as it is equal to the DDIM reference distribution. We assume (30) and (31) holds when $t=t_{0}\in\left\{1,2,\dots,t_{a}\right\}$, then when $t=t_{0}-1$, we have
[…]
$q_{\sigma}(x_{t_{0}-1}|x_{t_{0}},x_{0})$ is a Gaussian distribution according to the DDIM refernece distribution, and $q_{\sigma}(x_{t_{0}}|x_{t_{a}},x_{0})$ is a Gaussian distribution due to the inductive hypothesis. Then by the convolution law of two Gaussian distributions, (30) and (31) hold when $t=t_{0}-1$. By induction, Lemma 1 concludes. ∎
@@
Proof.
By Bayes’s formula,
$q_{\sigma}(x_{t-1}|x_{t},x_{t_{a}})=\frac{q_{\sigma}(x_{t}|x_{t-1},x_{t_{a}})q_{\sigma}(x_{t-1}|x_{t_{a}})}{q_{\sigma}(x_{t}|x_{t_{a}})}=\frac{q_{\sigma}(x_{t}|x_{t-1})q_{\sigma}(x_{t-1}|x_{t_{a}})}{q_{\sigma}(x_{t}|x_{t_{a}})}.$ (35)
Substituting the transition probability for $t\geq t_{a}$, we obtain the conclusion of Lemma 34. ∎
@@
A.2 Proof of Proposition 1
Proof.
Based on the above preliminaries, we derive the ELBO of $\log p_{\bm{\theta}}(x_{0},x_{t_{a}})$ as follows:
@@ A.3 Proof of Proposition 2
where the coefficient $g(t)$ is omitted as it can be scaled into the step size. Here we reach the result of Proposition 2 in the main text. ∎
A.4 Gradient truncation and remarks
In practice, we truncate the gradient through $\bm{\bm{\mu}_{\bm{\theta}}}$, i.e., we let
@@ C.1 Additional results
Here, we present the omitted experiments for image restoration tasks. The results for noise-free restoration tasks on ImageNet are shown in Table 5, the results for noise-free and noisy restoration tasks on CelebA are shown in Table 6 and 7, respectively, and the results for nonlinear high dynamic range (HDR) task are in Table 8. On the CelebA dataset, we further compare ProjDiff with more diffusion based rithms, including $\Pi$GDM [22], DMPS [52], Resample [34], and DiffPIR [53]. ProjDiff demonstrates highly competitive performance on all these tasks.
@@
C.2 Ablation study
Four trials for the phase retrieval task. It is noted in [23] that in the phase retrieval task, DPS requires sampling four times for each image and selecting the best result to achieve satisfactory performance. Therefore, we compare the performance of DPS and ProjDiff with four independent samplings. We also report the performance of the ER, HIO, and OSS algorithms with four samplings as baselines. The best image is selected as the sample with the highed PSNR metric. The results are shown in Table 11. Both DPS and ProjDiff show performance improvements with four trials, with ProjDiff in particular exhibiting much better performance, reaching a PSNR of 41.58dB in the noise-free case.
NFEs Method Phase Retrieval $\sigma=0$ Phase Retrieval $\sigma=0.1$
[…]
- ER-4trials 11.36 / 0.20 / 83.95 / 410.75 11.32 / 0.20 / 83.87 / 410.53
- HIO-4trials 12.60 / 0.26 / 79.78 / 336.94 12.62 / 0.26 / 79.88 / 333.90
- OSS-4trials 12.75 / 0.26 / 79.95 / 342.36 12.69 / 0.25 / 80.37 / 344.84
1000 DPS-4trials 19.36 / 0.45 / 48.48 / 104.70 17.25 / 0.38 / 51.83 / 82.10
1000 ProjDiff-4trials 41.58 / 0.94 / 5.49 / 7.19 28.11 / 0.76 / 28.43 / 53.38
The impact of the gradient truncation method used in ProjDiff. We compared ProjDiff with and without the gradient truncation method on the CelebA dataset, as shown in Table 12. ProjDiff without gradient truncation is denoted as ProjDiff-FG. The results indicates that there is some performance loss when using gradient truncation compared to not using it, while the efficiency increases by nearly three times. Considering that using gradient truncation has already achieved satisfactory performance, we accept the trade-off of some performance loss for the gain in efficiency.
Table 12: Ablation study of the gradient truncation method in ProjDiff on noise-free tasks on CelebA.
Time (s) Method Super-Resolution PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ Inpainting PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ Gaussian Deblurring PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
4.42 ProjDiff 32.57 / 0.90 / 16.08 / 33.74 36.84 / 0.96 / 6.66 / 12.27 45.57 / 0.99 / 2.99 / 2.05
11.13 ProjDiff-FG 33.05 / 0.91 / 16.87 / 36.39 37.30 / 0.96 / 7.12 / 11.65 45.98 / 0.99 / 2.72 / 1.94
Using gradient descent to approximate the projection operation for nonlinear tasks. We claim that for general nonlinear tasks, ProjDiff can be applied by using gradient descent to approximate the projection operation. Here we verify this method in the phase retrieval task (ProjDiff-GD in Table 13). The results indicate that approximating the projection operator of $x_{0}$ with respect to the observation equation by minimizing $||y-\mathcal{A}(x)||^{2}$ starting from $x_{0}$ can also achieve satisfactory performance, which suggests that ProjDiff has generalizability for general nonlinear inverse problems.
Table 13: Using gradient descent to approximate the projection operator.
NFEs Method Phase Retrieval $\sigma=0$ Phase Retrieval $\sigma=0.1$
[…]
1000 ProjDiff 33.39 / 0.74 / 20.42 / 35.91 23.84 / 0.60 / 38.55 / 76.20
1000 ProjDiff-GD 33.55 / 0.71 / 23.04 / 45.44 23.08 / 0.56 / 44.86 / 97.13
Sensitivity tests for the hyperparameters. We conduct the sensitivity tests for the step size $\eta_{1}$ and noise level $\sigma_{0}$ in ProjDiff on noisy super-resolution task on CelebA, as shown in Table 14. ’×$\times a$’ denotes that we perturb the input standard deviation of ProjDiff by multiplying $a$. The results indicate that ProjDiff exhibits a certain degree of robustness to the step size and noise level. This also demonstrates that retaining the equivalent noise level as an adjustable hyperparameter is feasible when it cannot be directly calculated.
Table 14: Ablation study of the hyperparameters in ProjDiff on noisy super-resolution task on CelebA.
PSNR $\uparrow$ LPIPS $\downarrow$
$\eta_{1}$\$\sigma_{0}$$\times 0.7$$\times 0.9$$\times 1.0$$\times 1.1$$\times 1.3$$\times 0.7$$\times 0.9$$\times 1.0$$\times 1.1$$\times 1.3$
0.1 29.92 30.16 30.19 30.18 30.11 21.96 21.71 22.25 22.70 23.29
0.2 29.69 29.94 29.96 29.94 29.85 22.48 20.92 21.23 21.60 22.12
0.4 29.16 29.47 29.49 29.47 29.37 24.25 21.04 20.86 21.06 21.51
0.8 28.45 28.85 28.89 28.87 28.75 27.67 22.82 21.85 21.63 21.74
1.6 27.69 28.10 28.14 28.09 27.88 32.23 26.62 24.88 24.07 23.69
@@ E.2 Hyperparameters for the compared algorithms
DDNM. DDNM uses the recommended parameter $\eta=0.85$ from [21]. For noisy tasks, we report the performance of the DDNM+ algorithm [21].
DPS. Following the settings in [23], the step size is set to $\xi_{t}=\xi/\left|\left|y-\mathcal{A}(\bm{\bm{\mu}_{\bm{\theta}}}(x_{t}))\right|\right|_{2}$. On ImageNet, $\xi$ is set to $1$ for super-resolution and inpainting tasks, and $0.4$ for Gaussian deblurring. Since [23] did not conduct experiments on CelebA, we conducted task-by-task parameter tuning for the learning rate in DPS and reported the results, ensuring a fair comparison. For phase retrieval and high dynamic range tasks, $\xi=0.4$.
@@
$\Pi$GDM. We carefully conducted hyperparameter searches for each task using the first eight images with the average PSNR as the metric. Except for Noisy Gaussian Deblur task, we avoided using other algorithms to initialize $\Pi$GDM to ensure a fair comparison. Specifically, we first use the forward transition to map the degraded image to the noise level corresponding to the 500th step, and then use 100 steps of $\Pi$GDM to solve the inverse problem. We find that this approach is much better than starting directly from white noise (i.e. 1000th step). For Noisy Gaussian Deblur task, $\Pi$GDM without initialization from other algorithms struggled to achieve satisfactory results. Therefore, we use DDNM+ to obtain the sample at the 500th step and then switch to $\Pi$GDM to complete the solution.
DMPS. The step size hyperparameter $\lambda$ was searched for each task using the first eight images with the average PSNR as the metric.
Resample. The weight hyperparameter $\gamma$ was searched for each task using the first eight images with the average PSNR as the metric. ReSample was originally designed for Latent Diffusion, thus in our experiments, we consider the encoder-decoder as identity mappings.
DiffPIR. The weight hyperparameter $\lambda$ was searched for each task using the first eight images with the average PSNR as the metric.
@@ F.3 Details for nonlinear observations
Appendix G Discussion about the MAP framework
ProjDiff shares similarities with the Maximum A Posteriori (MAP) estimation method. Theoretically, in noisy situations, it is only necessary to use a Gaussian likelihood, i.e., $p(y|x_{0})=\mathcal{N}(y;x_{0},\sigma^{2}I)$, to solve the inverse problems. However, ProjDiff introduces a new auxiliary variable, which leads to better experimental outcomes than the MAP framework. Here, we present some qualitative discussion towards this experimental result.
On one hand, if we are given perfectly accurate priors and likelihoods, and have sufficient computational capability, we could obtain the exact posterior and also its gradients. Within the MAP framework, this should lead to the ideal results. However, in practice, the priors provided by diffusion models are not entirely accurate, and the weight coefficients between the priors and the likelihood cannot always be perfectly set. Moreover, dealing with the priors in diffusion models relies on stochastic optimization methods. These factors imply that the MAP framework often fails to achieve the desired solution. In such cases, further exploring the information or capabilities within the diffusion model can provide substantial assistance in solving inverse problems and enhance performance. This is one of the reasons why ProjDiff may have a performance advantage over the original MAP framework.
On the other hand, MAP methods utilize gradient descent to leverage the information from the observation, while ProjDiff transforms this into a projection operation by introducing noisy auxiliary variables. This approach offers numerous advantages: there is no need to consider the step size of gradient descent (at least for the likelihood term); there is no need to consider the weight coefficients between the likelihood and prior terms; and it ensures consistency between noisy samples and observations (the role of this consistency has also been confirmed in the noise-free scenario). Moreover, the introduction of this auxiliary variable indicates that ProjDiff can actually be viewed as simultaneously recovering clean data and the noise added to the observations, which can yield more accurate results than merely characterizing the noise prior with a Gaussian distribution. These transformations are all thanks to ProjDiff’s utilization of both the clean prior and noisy prior modeled by diffusion models (i.e., the prior modeled by the diffusion model and its denoising capability).
Appendix H Algorithm blocks